# Two Optimizers Are Better Than One: LLM Catalyst Empowers Gradient-Based Optimization for Prompt Tuning

### v3

@@ 4.2 Main Results
Interpretation of prompts optimized by LLM. To further analyze the contribution of LLM-based optimizer in prompt optimization, we list the prompts contained in the instruction and generated by LLM in Figure 4. In round 1, the gradient-based optimizer tend to navigate around senseless prompt tokens, e.g., "beh", "ila", etc. This phenomenon is alleviated after leveraging LLM to infer for more meaningful prompts. In round 3, prompts with both interpretability and low loss values are obtained by the collaboration of gradient-based optimizer and LLM.
4.3 Ablation Study
Our method’s sensitivity to the choice of LLMs, prompt tuning baseline methods and amount of training samples are already shown in Table 1. We provide ablation results of three factors of our method in this section. The ablation experiments are based on CoOp baseline, using GPT-3.5 as optimizer. More ablation studies can be found in the Appendix.
Design of instruction. The instruction from which LLMs infer for new candidates is influential on the results. We empirically analyze the effect of each component in our instructions in Table 3. Task definition (TD) denotes raw instruction defining the task information. Manual prompt (MP) means LLMs are instructed with hand-crafted prompt template. Optimization trajectory (OT) denotes the intermediate results from the gradient optimizer is provided. The results in first line of Table 3 correspond to no LLM-based optimization, serving as a stronger baseline than CoOp. Random noise is added to perturb the prompts after each round of gradient-based optimization for the opportunity to escape from local optima. The ablation results show that hand-crafted template, providing prior knowledge of the prompt, and optimization trajectory, providing timely semantic landscape of current optimized prompts, are both important components for ideal performance.
@@ 4.3 Ablation Study
Rounds of alternating optimization. We analyze the effect of the alternating rounds $N$ of the two optimizer on the result. Table 5 indicate that the optimal round for each task varies. But more rounds involve more interactions with LLM, providing more candidates prompts. The average performance improves with more rounds generally. We choose 3 rounds as a proper value.
Iterations of gradient optimizer. We analyze the effect of the training iterations $m$ of gradient optimizer in every round. From Table 5, we can see that increasing training iterations leads to descending results. A longer training process by gradient-based optimizer may not necessarily benefit the final prompt tuning results. We find that more iterations of gradient-based optimization may results in candidate prompts with less semantic diversity, which is not good for proposing LLM to generate more promising candidate prompts. Thus, we use a small training iteration for better performance and efficiency of our optimization method.
@@ 5 Conclusion
In this paper, we proposed a combined optimization method that integrates LLM-based optimizer with conventional gradient-based optimization. Our method mitigates the inherent limitations of gradient-based optimization, such as entrapment in local optima, by leveraging the inductive reasoning capabilities of LLMs. By interleaving gradient-based optimization process with natural language interactions with LLMs, we introduced a higher-level guidance for optimization that effectively utilizes task descriptions and real-time optimization trajectories. We validated our combined optimization strategy through prompt tuning tasks, where the synergy between LLM-based optimizer and gradient-based optimizer has consistently demonstrated improved performance over competitive baselines. These results underscore the complementary effect of LLM-based optimizer and conventional gradient-based optimization. Our contributions inspire further exploration of the advantages of LLM-based optimization over existing algorithms, paving the way for more effective integration of LLM-based inference into conventional optimization workflows.
Limitations. Our proposed method can not be effectively employed for adapter-based fine-tuning methods. Feasible solution of expressing optimization status of adapter module to LLMs needs further design. We leave the application of the proposed method to more wider range of optimization problems (e.g., adapters, LoRA) and algorithms (e.g., reinforce learning) as future work.
@@ A.1 More Experimental Details
The design of instruction for Llama2-7B is different from GPT-3.5 and GPT-4, since we notice that the instruction following ability of Llama2-7B is weaker. It is more likely to produce unexpected output. Even though we emphasized desired way of responding to our query, the responses from Llama2-7B still need proper post-processing to obtain the clean returned prompts.
System: You are a helpful, respectful and honest assistant capable of proposing new prompts for users. User: Propose new prompts for user. Reply with only the proposed short template, do not reply the loss and accuracy. Keep every template under 8 words. Generate 3 templates that potentially have better image recognition performance. I have a list of text templates with their corresponding loss values and accuracy. They are used for image classification with CLIP model. The templates are arranged in descending order based on their loss value on training samples, where lower loss indicates better quality. (Insert optimized prompts as optimization trajectories here.)
@@
Detailed hyperparameter settings. The backbone models used by CoOp and TCP are ResNet50 and ViT-B/16 respectively. The prompt length is set as 4 for both CoOp and TCP. The training hyperparameters, such as epochs and learning rate, are remained the same as the original methods. The number of training iterations $M$ for Algorithm 1 is equal to the training iterations of the original methods. We set the number of rounds $N$ as 3 and the iteration for the gradient optimizer $m$ is set as 10 for CoOp, 30 for TCP. The prompt length for NLU tasks is set as 8.
@@ A.2 More Ablation Study
Distance function used for token space projection. The token space projection operator in Eqn. 3 uses L2 distance to find the nearest discrete tokens for continuous prompt embeddings. We also tried to using cosine similarity as distance function. The results are provided in Table 6
@@
Length of the prompt. We use a default prompt length 4 for our experiments. We provide the result of our method with longer prompt in Table 7.
@@
Prompt Length EuroSAT DTD Oxford_Flowers
4 71.39 56.70 86.64
8 70.52 56.38 86.90

### v4

@@ 4.3 Ablation Study
Design of instruction. The instruction from which LLMs infer for new candidates influences the results. We empirically analyze the effect of each component in our instructions in Table 4. Task definition (TD) denotes raw instruction defining the task information. Manual prompt (MP) means LLMs are instructed with hand-crafted prompt templates. Optimization trajectory (OT) denotes the intermediate results from the gradient optimizer provided. The results on the first line of Table 4 correspond to no LLM-based optimization, serving as a baseline. The ablation results show that hand-crafted templates, providing prior knowledge of the prompt, and optimization trajectory, providing a timely semantic landscape of currently optimizing prompts, are both important components for ideal performance.
@@
Rounds of alternating optimization. We analyze the effect of the alternating rounds $N$ of the two optimizers on the result. Table 6 indicates that the optimal round for each task varies. But more rounds involve more interactions with LLM, providing more candidates prompts. The average performance improves with more rounds generally. We choose 3 rounds as a proper value.
The timing of interaction between two optimizers. We explored how the timing of interactions between the LLM optimizer and the gradient optimizer affects optimization results, maintaining a constant total number of gradient optimization iterations (i.e., keeping $m\times N+M$ iterations constant). In reference to Table 6, smaller values of $m$ indicate that the LLM optimizer is involved early in the optimization process, whereas larger values of $m$ indicate that the LLM optimizer is introduced during the latter stages of the optimization process. We find that larger $m$ may result in candidate prompts in the gradient trajectory with less semantic diversity, which is less effective for proposing LLM to generate more promising candidate prompts. Furthermore, larger $m$ means smaller $M$ for the last round of gradient optimization, which may lead to insufficient convergence of the algorithm, degrades performance. Thus, we employ a smaller number of training iterations to enable the LLM optimizer to offer a rich variety of candidate prompts during the initial stages of optimization.
@@ 5 Conclusion
This paper proposes a collaborative optimization method combining the conventional gradient-based optimizer and inferential LLM-based optimizer. By alternating between the gradient-based and LLM-based optimization process, we combine the local carefulness of gradient- based optimizer and diverse semantic exploration of LLM-based optimizer. LLM-based optimizer mitigates the inherent limitations of gradient-based optimization, such as entrapment in local optima, by inferring high-level guidance from task descriptions and real-time optimization trajectories. We validated our combined optimization method through prompt tuning tasks, where the synergy between LLM-based optimizer and gradient-based optimizer has consistently demonstrated improved performance over competitive baselines. These results underscore the complementary effect of LLM-based optimizer and conventional gradient-based optimization. Our contributions inspire further exploration of the advantages of LLM-based optimization over existing algorithms, paving the way for more effective integration of LLM-based inference into conventional optimization workflows.
Limitations. Our proposed optimization method can not be directly employed for adapter-based or LoRA-based fine-tuning methods. A feasible solution for handling higher dimensional parameters in LLM-based optimization needs to be designed. We leave the application of the proposed optimization framework to broader range of optimization problems (e.g., adapters, LoRA) and algorithms (e.g., reinforce learning) as future work.
@@ A.1 More Experimental Details
The design of instruction for Llama2-7B is different from GPT-3.5 and GPT-4 since we notice that the instruction following ability of Llama2-7B is weaker. It is more likely to produce unexpected output. Even though we emphasized the desired way of responding to our query, the responses from Llama2-7B still need proper post-processing to obtain the clean returned prompts.
System: You are a helpful, respectful, and honest assistant capable of proposing new prompts for users. User: Propose new prompts for user. Reply with only the proposed short template, do not reply the loss and accuracy. Keep every template under 8 words. Generate 3 templates that potentially have better image recognition performance. I have a list of text templates with their corresponding loss values and accuracy. They are used for image classification with CLIP model. The templates are arranged in descending order based on their loss value on training samples, where lower loss indicates better quality. (Insert optimized prompts as optimization trajectories here.)
@@
Detailed hyperparameter settings. The backbone models used by CoOp and TCP are ResNet50 and ViT-B/16, respectively. The prompt length is set as 4 for both CoOp and TCP. The training hyperparameters, such as epochs and learning rate, remained the same as the original methods. The number of training iterations $M$ for Algorithm 1 equals the training iterations of the original methods. We set the number of rounds $N$ as 3, and the iteration for the gradient optimizer $m$ is set as 10 for CoOp and 30 for TCP. The prompt length for NLU tasks is set as 8. The experiments are conducted on a V100 GPU. The specific versions of the API we are utilizing are “gpt-3.5-turbo-1106" for “GPT-3.5" and “gpt-4-1106-preview" for “GPT-4".
@@ A.2 More Ablation Study
Distance function used for token space projection. The token space projection operator in Eqn. 3 uses L2 distance to find the nearest discrete tokens for continuous prompt embeddings. We also tried to use cosine similarity as a distance function. The results are provided in Table 7
@@
Length of the prompt. We use a default prompt length of 4 for our experiments. We provide the result of our method with a longer prompt in Table 8.
@@
Methods EuroSAT DTD Oxford_Flowers
Gradient-based Search (length 4) 69.14 54.96 85.04
Ours (length 4) 71.39 56.70 86.64
Gradient-based Search (length 8) 69.36 55.10 85.48
Ours (length 8) 70.52 56.38 86.90
Gradient-based Search (length 16) 70.55 54.93 85.01
Ours (length 16) 71.54 56.24 86.45